# Preliminaries on Pre-Hilbert Structures on Polynomial Spaces and Associated Laplacians

### v1

@@ 5.2. Resolvent-based distance
The map $d_{\mathrm{res}}$ defines a metric on the set of polynomial Hilbert geometries modulo unitary equivalence.
@@
5.3. Finite-degree truncations and stability theorem
Let $\mathcal{P}_{\leq N}(\Omega)$ denote the space of polynomials of degree at most $N$, and let $P_{N}:H\to\mathcal{P}_{\leq N}(\Omega)$ be the orthogonal projection.
@@ 5.3. Finite-degree truncations and stability theorem
The truncated Laplacian of order $N$ associated with $\mathfrak{G}$ is the finite-rank operator
$\Delta_{\mathfrak{G}}^{(N)}:=P_{N}\Delta_{\mathfrak{G}}P_{N}.$
This operator admits a matrix representation in any orthonormal polynomial basis adapted to the degree filtration. Throughout, for each $N$ we denote by $P_{N}$ the orthogonal projector onto the closed subspace $\mathcal{P}_{\leq N}$ generated by polynomials of degree $\leq N$, and we define the truncated resolvent by
$(1+\Delta^{(N)})^{-1}:=P_{N}(1+\Delta)^{-1}P_{N}\quad\text{as an operator on }H_{\leq N}.$
@@ 6.3. Stability of polynomial kernels
Corollary 6.5 (Kernel stability).
Under the assumptions of Theorem 6.1,
$\|K_{N}^{(1)}-K_{N}^{(2)}\|_{L^{\infty}(\Omega\times\Omega)}\leq C_{N}\,\varepsilon.$
@@
The kernel is the integral kernel of the orthogonal projector $P_{N}$. Operator norm control of $P_{N}^{(1)}-P_{N}^{(2)}$ implies uniform control of the kernels on compact sets. ∎
6.4. Interpretation
Theorems 6.1 and 6.3 show that the norm-resolvent topology on polynomial Hilbert geometries induces a robust notion of stability for:
  • polynomial subspaces;
  • Gram–Schmidt orthogonal bases;
  • reproducing kernels.
This justifies the use of resolvent-based distances as a quantitative tool to compare polynomial orthogonalization procedures.
@@ 7.1. Classical orthogonal polynomials on an interval
$\mathfrak{G}_{w}=(\mathcal{P}(I),\langle\cdot,\cdot\rangle_{w},\partial_{x}),\qquad\langle f,gImgle_{w}=\int_{I}fg\,w\,dx.$
@@ 7.2. Sobolev orthogonal polynomials on an interval
$\langle f,gImgle_{S}=\sum_{k=0}^{s}\lambda_{k}\int_{I}f^{(k)}(x)g^{(k)}(x)\,dx,\qquad\lambda_{0}>0,\ \lambda_{k}\geq 0.$
@@ 7.4. An explicit model on $S^{1}$: weighted $L^{2}$ versus Sobolev orthogonalization
the standard Fourier orthonormal basis of $L^{2}$. Let $D=\partial_{\theta}$ defined on trigonometric polynomials (dense in $L^{2}$), so that
@@
$\langle f,gImgle_{w}:=\int_{0}^{2\pi}f(\theta)\overline{g(\theta)}\,w(\theta)\,\frac{d\theta}{2\pi}=\langle M_{w}f,gImgle_{L^{2}},$
@@
(7.1) $\langle f,gImgle_{w,\lambda}:=\int f\overline{g}\,w\,\frac{d\theta}{2\pi}\;+\;\lambda\int f^{\prime}(\theta)\,\overline{g^{\prime}(\theta)}\,\frac{d\theta}{2\pi}.$
@@
(7.2) $\langle f,gImgle_{w,\lambda}=\langle(M_{w}+\lambda\Delta_{0})f,gImgle_{L^{2}}.$
@@ 7.4.1. Adjoint of $D$ and Laplacian in the Sobolev geometry
$\langle Df,gImgle_{w,\lambda}=\langle A_{w,\lambda}Df,gImgle_{L^{2}}=\langle Df,A_{w,\lambda}gImgle_{L^{2}}.$
@@
$\langle Df,gImgle_{w,\lambda}=\langle f,D^{*}(A_{w,\lambda}g)Imgle_{L^{2}}=\langle f,A_{w,\lambda}\,A_{w,\lambda}^{-1}D^{*}A_{w,\lambda}gImgle_{L^{2}}=\langle f,A_{w,\lambda}^{-1}D^{*}A_{w,\lambda}gImgle_{w,\lambda}.$
@@ 7.4.5. A Fourier-based proof of the resolvent bound
$\langle\Delta_{w,0}u,uImgle_{w}=\langle Du,DuImgle_{w}=\int_{0}^{2\pi}|u^{\prime}(\theta)|^{2}\,w(\theta)\,\frac{d\theta}{2\pi}\geq w_{-}\,\|u^{\prime}\|_{L^{2}}^{2}.$
@@
(7.12) $\langle(1+\Delta_{w,0})u,uImgle_{w}\geq\|u\|_{L^{2}(w)}^{2}+w_{-}\,\|u^{\prime}\|_{L^{2}}^{2}.$
@@
$\|u\|_{L^{2}(w)}^{2}+w_{-}\,\|u^{\prime}\|_{L^{2}}^{2}\leq\langle f,uImgle_{w}\leq\|f\|_{L^{2}(w)}\|u\|_{L^{2}(w)},$
@@ 7.4.6. Auxiliary Fourier lemmas on $S^{1}$
$\langle A_{w,\lambda}u,uImgle_{L^{2}}\geq w_{-}\,\|u\|_{L^{2}}^{2}.$
@@
From the energy identity $\langle\Delta_{w,0}u,uImgle_{w}=\|u^{\prime}\|_{L^{2}(w)}^{2}$, we obtain the $H^{1}$ bounds
@@
and let $Q_{N}$ be the orthogonal projector onto $\mathcal{T}_{\leq N}$. We view $\mathcal{T}_{\leq N}$ as a finite-dimensional Hilbert space with inner product $\langle f,gImgle_{w}=\int f\overline{g}\,w\,d\theta/2\pi$ (equivalent to the $L^{2}$ inner product).
@@
8. Limit polynomial geometries and resolvent rigidity
8.1. Resolvent distance as a metric on polynomial geometries
The constructions and results of the previous sections show that the resolvent distance
$d_{\mathrm{res}}(\mathfrak{G}_{1},\mathfrak{G}_{2})=\big\|(1+\Delta_{\mathfrak{G}_{1}})^{-1}-(1+\Delta_{\mathfrak{G}_{2}})^{-1}\big\|$
defines a natural metric on polynomial Hilbert geometries, modulo unitary equivalence.
Unlike moment-based or weak topologies on measures, the resolvent distance is intrinsically operatorial: it compares the geometric content of the orthogonalization procedure rather than its probabilistic representation. In particular, convergence in resolvent distance implies uniform stability of all finite-degree polynomial subspaces, Gram–Schmidt bases and reproducing kernels (Theorems 6.1 and 6.3).
8.2. Limit geometries beyond measures
A striking consequence of this approach is that the space of polynomial geometries is strictly larger than the space of measures. While classical orthogonal polynomials are generated by measures and correspond to Sturm–Liouville operators diagonalized by polynomial bases, Sobolev and fractional Sobolev inner products lead to Laplacians with banded or almost-banded matrix representations.
The examples on $S^{1}$ and on thin annuli illustrate that sequences of polynomial geometries may converge in resolvent distance to limit objects which are not associated with any finite Radon measure. Such limits are nevertheless perfectly well defined at the level of Laplacians and orthogonal polynomial structures. This provides a natural framework to study degenerations, regularizations and geometric limits of orthogonal polynomials.
8.3. Rigidity and flexibility
In the classical cases (Jacobi, Laguerre, Hermite, OPUC), the associated Laplacians are exactly diagonalizable in the orthogonal polynomial basis, reflecting a maximal rigidity. Any perturbation preserving diagonalizability forces the geometry to remain within the same classical family.
By contrast, Sobolev-type geometries exhibit controlled flexibility: the Laplacian matrices are banded or almost banded, and the resolvent distance provides a quantitative measure of deviation from the classical regime. Thin annulus limits show how higher-dimensional geometries collapse onto lower-dimensional ones while preserving resolvent control.
8.4. Perspectives
The resolvent-based approach developed here opens several directions:
  • the study of polynomial geometries induced by nonlocal or pseudo-differential energies;
  • quantitative comparison of orthogonalization procedures arising in numerical analysis and learning theory;
  • extensions to multivariate and noncommutative polynomial settings.
More generally, it suggests viewing orthogonal polynomials as manifestations of an underlying operator geometry, rather than as objects tied to a specific measure.

### v2

@@ 5.2. Resolvent-based distance
Let $H$ be a Hilbert space and let $\Delta_{1},\Delta_{2}$ be nonnegative self-adjoint operators on $H$. Define
$d_{\mathrm{res}}(\Delta_{1},\Delta_{2}):=\big\|(1+\Delta_{1})^{-1}-(1+\Delta_{2})^{-1}\big\|_{\mathcal{B}(H)}.$
Then $d_{\mathrm{res}}$ is a metric on the set of nonnegative self-adjoint operators on $H$. Moreover, for every unitary $U\in\mathcal{U}(H)$,
$d_{\mathrm{res}}(U\Delta_{1}U^{-1},\,U\Delta_{2}U^{-1})=d_{\mathrm{res}}(\Delta_{1},\Delta_{2}).$
Proof.
Step 0: boundedness of resolvents. Since each $\Delta_{i}$ is self-adjoint and nonnegative, its spectrum satisfies $\sigma(\Delta_{i})\subset[0,\infty)$. Hence $1\notin\sigma(\Delta_{i})$ and the resolvent
$R_{i}:=(1+\Delta_{i})^{-1}$
exists as a bounded operator on $H$. Moreover, by the spectral theorem,
$\|R_{i}\|_{\mathcal{B}(H)}=\sup_{\lambda\in\sigma(\Delta_{i})}\frac{1}{1+\lambda}\leq 1.$
Therefore $d_{\mathrm{res}}(\Delta_{1},\Delta_{2})=\|R_{1}-R_{2}\|$ is well-defined and finite.
Step 1: positivity and symmetry. By definition, $d_{\mathrm{res}}(\Delta_{1},\Delta_{2})$ is the norm of a bounded operator, hence $d_{\mathrm{res}}(\Delta_{1},\Delta_{2})\geq 0$. Also,
$d_{\mathrm{res}}(\Delta_{1},\Delta_{2})=\|R_{1}-R_{2}\|=\|-(R_{2}-R_{1})\|=d_{\mathrm{res}}(\Delta_{2},\Delta_{1}),$
so $d_{\mathrm{res}}$ is symmetric.
Step 2: triangle inequality. Let $\Delta_{1},\Delta_{2},\Delta_{3}$ be nonnegative self-adjoint on $H$ and set $R_{i}=(1+\Delta_{i})^{-1}$. Then
$R_{1}-R_{3}=(R_{1}-R_{2})+(R_{2}-R_{3}),$
and by the triangle inequality in the operator norm,
$d_{\mathrm{res}}(\Delta_{1},\Delta_{3})=\|R_{1}-R_{3}\|\leq\|R_{1}-R_{2}\|+\|R_{2}-R_{3}\|=d_{\mathrm{res}}(\Delta_{1},\Delta_{2})+d_{\mathrm{res}}(\Delta_{2},\Delta_{3}).$
Step 3: separation (identity of indiscernibles). Assume $d_{\mathrm{res}}(\Delta_{1},\Delta_{2})=0$. Then $\|R_{1}-R_{2}\|=0$, hence $R_{1}=R_{2}$ as bounded operators. We claim that this implies $\Delta_{1}=\Delta_{2}$.
Indeed, since $\Delta_{i}$ is self-adjoint, the operator $1+\Delta_{i}$ is self-adjoint with domain $\mathrm{Dom}(\Delta_{i})$, and it is bijective from $\mathrm{Dom}(\Delta_{i})$ onto $H$, with bounded inverse $R_{i}$. In particular,
$(1+\Delta_{i})R_{i}=I\quad\text{on }H,\qquad R_{i}(1+\Delta_{i})=I\quad\text{on }\mathrm{Dom}(\Delta_{i}).$
Since $R_{1}=R_{2}=:R$, we have $\mathrm{Ran}(R)=\mathrm{Dom}(\Delta_{1})=\mathrm{Dom}(\Delta_{2})$. Moreover, on this common domain,
$(1+\Delta_{1})u=R^{-1}u=(1+\Delta_{2})u.$
Hence $\Delta_{1}u=\Delta_{2}u$ for all $u\in\mathrm{Dom}(\Delta_{1})=\mathrm{Dom}(\Delta_{2})$, so $\Delta_{1}=\Delta_{2}$. Therefore $d_{\mathrm{res}}(\Delta_{1},\Delta_{2})=0$ implies $\Delta_{1}=\Delta_{2}$.
Conversely, if $\Delta_{1}=\Delta_{2}$, then $R_{1}=R_{2}$ and $d_{\mathrm{res}}(\Delta_{1},\Delta_{2})=0$. This proves separation.
Step 4: unitary invariance. Let $U$ be unitary on $H$. If $\Delta$ is self-adjoint and nonnegative, then so is $\widetilde{\Delta}:=U\Delta U^{-1}$, with $\sigma(\widetilde{\Delta})=\sigma(\Delta)$. Moreover,
$1+\widetilde{\Delta}=U(1+\Delta)U^{-1},$
hence by taking inverses on both sides,
$(1+\widetilde{\Delta})^{-1}=U(1+\Delta)^{-1}U^{-1}.$
Applying this to $\Delta_{1},\Delta_{2}$ gives
$d_{\mathrm{res}}(U\Delta_{1}U^{-1},\,U\Delta_{2}U^{-1})$ $=\big\|U(1+\Delta_{1})^{-1}U^{-1}-U(1+\Delta_{2})^{-1}U^{-1}\big\|$
$=\big\|U\big((1+\Delta_{1})^{-1}-(1+\Delta_{2})^{-1}\big)U^{-1}\big\|.$
Since conjugation by a unitary preserves the operator norm, $\|UTU^{-1}\|=\|T\|$, we obtain
$d_{\mathrm{res}}(U\Delta_{1}U^{-1},\,U\Delta_{2}U^{-1})=d_{\mathrm{res}}(\Delta_{1},\Delta_{2}).$
This completes the proof that $d_{\mathrm{res}}$ is a metric and is unitarily invariant. ∎
@@
Remark 5.7.
When comparing two polynomial Hilbert geometries, we will always assume that their Laplacians are realized on the same Hilbert space (or transported to a common Hilbert space through a fixed unitary identification). In that case $d_{\mathrm{res}}$ is the usual norm-resolvent distance between the transported Laplacians. We do not claim that it separates geometries inducing the same Laplacian.
5.3. Finite-degree truncations
Let $\mathcal{P}_{\leq N}(\Omega)$ denote the space of polynomials of degree at most $N$, and let $H_{\leq N}\subset H$ be its (finite-dimensional) image in $H$. We denote by $P_{N}:H\to H_{\leq N}$ the orthogonal projector.
@@ 5.3. Finite-degree truncations
The truncated Laplacian of order $N$ associated with $\mathfrak{G}$ is the operator on $H_{\leq N}$
$\Delta_{\mathfrak{G}}^{(N)}:=P_{N}\,\Delta_{\mathfrak{G}}\big|_{H_{\leq N}}:H_{\leq N}\to H_{\leq N}.$
Definition 5.9 (Compressed resolvent).
We also consider the compressed resolvent on $H_{\leq N}$:
$R_{\mathfrak{G}}^{\langle N\rangle}:=P_{N}(1+\Delta_{\mathfrak{G}})^{-1}P_{N}\in\mathcal{B}(H_{\leq N}).$
Remark 5.10.
In general one has $R_{\mathfrak{G}}^{\langle N\rangle}\neq(I+\Delta_{\mathfrak{G}}^{(N)})^{-1}$. They coincide when $H_{\leq N}$ is invariant under $\Delta_{\mathfrak{G}}$ (e.g. if the derivation preserves the degree filtration).
@@ 6.3. Stability of polynomial kernels
Corollary 6.5 (Kernel stability on compacts).
Assume $\Omega$ is a metric space and fix a compact set $K\subset\Omega$. For $i=1,2$, let
$K_{N}^{(i)}(x,y)=\sum_{k=0}^{N}p_{k}^{(i)}(x)\overline{p_{k}^{(i)}(y)}$
be the polynomial reproducing kernel of $H_{\leq N}$ (in any orthonormal basis). Then there exists a constant $C_{N,K}>0$ such that
$\sup_{(x,y)\in K\times K}\big|K_{N}^{(1)}(x,y)-K_{N}^{(2)}(x,y)\big|\leq C_{N,K}\,\|P_{N}^{(1)}-P_{N}^{(2)}\|_{\mathcal{B}(H)}.$
In particular, under the assumptions of Theorem 6.1, one has
$\sup_{K\times K}|K_{N}^{(1)}-K_{N}^{(2)}|\leq C^{\prime}_{N,K}\,\varepsilon.$
@@
Let $V_{i}:=\mathrm{Ran}(P_{N}^{(i)})\subset H$ (finite-dimensional), and set
$V:=V_{1}+V_{2}\subset H.$
All vectors in $V$ are (restrictions of) polynomial functions on $\Omega$, hence are continuous on $K$. For each $x\in K$, consider the conjugate evaluation functional on $V$,
$\Lambda_{x}:V\to\mathbb{C},\qquad\Lambda_{x}(f):=\overline{f(x)}.$
Since $V$ is finite-dimensional, $\Lambda_{x}$ is continuous for the $H$-norm. By the Riesz representation theorem (for continuous conjugate-linear functionals under our convention), there exists a unique vector $g_{x}\in V$ such that
(6.1) $\Lambda_{x}(f)=\langle f,g_{x}\rangle_{H}\qquad\forall f\in V.$
Step 1: reproducing vectors in $V_{i}$. For $i\in\{1,2\}$ define
$k_{x}^{(i)}:=P_{N}^{(i)}g_{x}\in V_{i}.$
Then for every $f\in V_{i}$ we have, using $P_{N}^{(i)}f=f$ and orthogonality of $P_{N}^{(i)}$,
$\overline{f(x)}=\Lambda_{x}(f)=\langle f,g_{x}\rangle_{H}=\langle P_{N}^{(i)}f,g_{x}\rangle_{H}=\langle f,P_{N}^{(i)}g_{x}\rangle_{H}=\langle f,k_{x}^{(i)}\rangle_{H}.$
Thus $k_{x}^{(i)}$ is the Riesz representer of $\Lambda_{x}$ restricted to $V_{i}$.
Step 2: identification with the kernel. Let $\{p_{0}^{(i)},\dots,p_{N}^{(i)}\}$ be any orthonormal basis of $V_{i}$ in $H$. Expanding $k_{x}^{(i)}$ in this basis and using $\langle p_{j}^{(i)},k_{x}^{(i)}\rangle_{H}=\overline{p_{j}^{(i)}(x)}$, we obtain
(6.2) $k_{x}^{(i)}=\sum_{j=0}^{N}\overline{p_{j}^{(i)}(x)}\,p_{j}^{(i)}.$
Therefore, for $x,y\in K$,
$\langle k_{y}^{(i)},k_{x}^{(i)}\rangle_{H}=\sum_{j=0}^{N}p_{j}^{(i)}(y)\,\overline{p_{j}^{(i)}(x)}=K_{N}^{(i)}(y,x).$
In particular,
(6.3) $\big|K_{N}^{(1)}(x,y)-K_{N}^{(2)}(x,y)\big|=\big|K_{N}^{(1)}(y,x)-K_{N}^{(2)}(y,x)\big|=\big|\langle k_{y}^{(1)},k_{x}^{(1)}\rangle_{H}-\langle k_{y}^{(2)},k_{x}^{(2)}\rangle_{H}\big|.$
Step 3: estimate by $\|P_{N}^{(1)}-P_{N}^{(2)}\|$. Using $k_{x}^{(i)}=P_{N}^{(i)}g_{x}$ and adding/subtracting,
$\langle P_{1}g_{y},P_{1}g_{x}\rangle_{H}-\langle P_{2}g_{y},P_{2}g_{x}\rangle_{H}=\langle(P_{1}-P_{2})g_{y},P_{1}g_{x}\rangle_{H}+\langle P_{2}g_{y},(P_{1}-P_{2})g_{x}\rangle_{H},$
where we wrote $P_{i}:=P_{N}^{(i)}$ for brevity. Hence, since $\|P_{i}\|=1$,
$\big|\langle k_{y}^{(1)},k_{x}^{(1)}\rangle_{H}-\langle k_{y}^{(2)},k_{x}^{(2)}\rangle_{H}\big|\leq\|P_{1}-P_{2}\|_{\mathcal{B}(H)}\big(\|g_{y}\|_{H}\,\|g_{x}\|_{H}+\|g_{y}\|_{H}\,\|g_{x}\|_{H}\big)=2\|P_{1}-P_{2}\|\,\|g_{x}\|\,\|g_{y}\|.$
Taking the supremum over $(x,y)\in K\times K$ and using (6.3), we obtain
(6.4) $\sup_{(x,y)\in K\times K}\big|K_{N}^{(1)}(x,y)-K_{N}^{(2)}(x,y)\big|\leq 2\Big(\sup_{x\in K}\|g_{x}\|_{H}\Big)^{2}\,\|P_{N}^{(1)}-P_{N}^{(2)}\|_{\mathcal{B}(H)}.$
Step 4: finiteness of the constant on compacts. It remains to check that
$M_{N,K}:=\sup_{x\in K}\|g_{x}\|_{H}<\infty.$
Choose a basis $(\phi_{1},\dots,\phi_{m})$ of $V$ and let $G\in\mathbb{C}^{m\times m}$ be its Gram matrix $G_{\alpha\beta}:=\langle\phi_{\beta},\phi_{\alpha}\rangle_{H}$ (invertible since it is positive definite on $V$). Writing $g_{x}=\sum_{\beta}c_{\beta}(x)\phi_{\beta}$ and using (6.1) with $f=\phi_{\alpha}$ gives
$\overline{\phi_{\alpha}(x)}=\langle\phi_{\alpha},g_{x}\rangle_{H}=\sum_{\beta}c_{\beta}(x)\langle\phi_{\alpha},\phi_{\beta}\rangle_{H}=\sum_{\beta}G_{\alpha\beta}\,c_{\beta}(x),$
hence $c(x)=G^{-1}\overline{\Phi(x)}$ with $\Phi(x)=(\phi_{1}(x),\dots,\phi_{m}(x))^{\top}$. Since each $\phi_{\alpha}$ is continuous on $K$, the map $x\mapsto\Phi(x)$ is continuous on the compact set $K$, so $\sup_{x\in K}\|\Phi(x)\|<\infty$. Therefore $\sup_{x\in K}\|c(x)\|<\infty$, and hence $\sup_{x\in K}\|g_{x}\|_{H}<\infty$. This proves $M_{N,K}<\infty$ and thus (6.4).
Setting $C_{N,K}:=2M_{N,K}^{2}$ yields the first inequality in the corollary.
Final step. Under the assumptions of Theorem 6.1, we have $\|P_{N}^{(1)}-P_{N}^{(2)}\|_{\mathcal{B}(H)}\leq C_{N}\,\varepsilon$, hence
$\sup_{K\times K}|K_{N}^{(1)}-K_{N}^{(2)}|\leq(C_{N,K}C_{N})\,\varepsilon,$
so one may take $C^{\prime}_{N,K}:=C_{N,K}C_{N}$. ∎
@@ 7.1. Classical orthogonal polynomials on an interval
$\mathfrak{G}_{w}=(\mathcal{P}(I),\langle\cdot,\cdot\rangle_{w},\partial_{x}),\qquad\langle f,g\rangle_{w}=\int_{I}fg\,w\,dx.$
@@ 7.2. Sobolev orthogonal polynomials on an interval
$\langle f,g\rangle_{S}=\sum_{k=0}^{s}\lambda_{k}\int_{I}f^{(k)}(x)g^{(k)}(x)\,dx,\qquad\lambda_{0}>0,\ \lambda_{k}\geq 0.$
@@ 7.4. An explicit model on $S^{1}$: weighted $L^{2}$ versus Sobolev orthogonalization
the standard Fourier orthonormal basis of $L^{2}$. Trigonometric polynomials appear here via the parametrization $t\mapsto e^{it}:$ for any polynomial $P\in\mathbb{C}[X,Y]$ the function $t\mapsto P(cost,sint)$ is a trigonometric polynomial.
Let $D=\partial_{\theta}$ defined on trigonometric polynomials (dense in $L^{2}$), so that
@@
$\langle f,g\rangle_{w}:=\int_{0}^{2\pi}f(\theta)\overline{g(\theta)}\,w(\theta)\,\frac{d\theta}{2\pi}=\langle M_{w}f,g\rangle_{L^{2}},$
@@
(7.1) $\langle f,g\rangle_{w,\lambda}:=\int f\overline{g}\,w\,\frac{d\theta}{2\pi}\;+\;\lambda\int f^{\prime}(\theta)\,\overline{g^{\prime}(\theta)}\,\frac{d\theta}{2\pi}.$
@@
(7.2) $\langle f,g\rangle_{w,\lambda}=\langle(M_{w}+\lambda\Delta_{0})f,g\rangle_{L^{2}}.$
@@ 7.4.1. Adjoint of $D$ and Laplacian in the Sobolev geometry
$\langle Df,g\rangle_{w,\lambda}=\langle A_{w,\lambda}Df,g\rangle_{L^{2}}=\langle Df,A_{w,\lambda}g\rangle_{L^{2}}.$
@@
$\langle Df,g\rangle_{w,\lambda}=\langle f,D^{*}(A_{w,\lambda}g)\rangle_{L^{2}}=\langle f,A_{w,\lambda}\,A_{w,\lambda}^{-1}D^{*}A_{w,\lambda}g\rangle_{L^{2}}=\langle f,A_{w,\lambda}^{-1}D^{*}A_{w,\lambda}g\rangle_{w,\lambda}.$
@@ 7.4.5. A Fourier-based proof of the resolvent bound
$\langle\Delta_{w,0}u,u\rangle_{w}=\langle Du,Du\rangle_{w}=\int_{0}^{2\pi}|u^{\prime}(\theta)|^{2}\,w(\theta)\,\frac{d\theta}{2\pi}\geq w_{-}\,\|u^{\prime}\|_{L^{2}}^{2}.$
@@
(7.12) $\langle(1+\Delta_{w,0})u,u\rangle_{w}\geq\|u\|_{L^{2}(w)}^{2}+w_{-}\,\|u^{\prime}\|_{L^{2}}^{2}.$
@@
$\|u\|_{L^{2}(w)}^{2}+w_{-}\,\|u^{\prime}\|_{L^{2}}^{2}\leq\langle f,u\rangle_{w}\leq\|f\|_{L^{2}(w)}\|u\|_{L^{2}(w)},$
@@ 7.4.6. Auxiliary Fourier lemmas on $S^{1}$
$\langle A_{w,\lambda}u,u\rangle_{L^{2}}\geq w_{-}\,\|u\|_{L^{2}}^{2}.$
@@
From the energy identity $\langle\Delta_{w,0}u,u\rangle_{w}=\|u^{\prime}\|_{L^{2}(w)}^{2}$, we obtain the $H^{1}$ bounds
@@
and let $Q_{N}$ be the orthogonal projector onto $\mathcal{T}_{\leq N}$. We view $\mathcal{T}_{\leq N}$ as a finite-dimensional Hilbert space with inner product $\langle f,g\rangle_{w}=\int f\overline{g}\,w\,d\theta/2\pi$ (equivalent to the $L^{2}$ inner product).